OpenStudy (anonymous):

Find the inverse of the function. f(x)=3squareroot x/8-4 i keep getting f−1(x)=512x3+6144x2+24576x+32768

OpenStudy (anonymous):

when these are the answer choices f-1(x) = [8(x + 4)]3 f-1(x) = 8(x + 4)3 f-1(x) = 8(x3 + 4) f-1(x) = 24(x + 4)

OpenStudy (anonymous):

Can you draw the function?

OpenStudy (anonymous):

Because it's not very clear. :)

OpenStudy (jack1):

hi... is this your function? \(\Large f(x)=3 \times \sqrt{ \frac x{8-4}}\)

OpenStudy (anonymous):

no

OpenStudy (jack1):

is it this? \(\Large f(x)=3 \times \sqrt{ \frac x8-4}\)

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x/8}-4\]

OpenStudy (anonymous):

Step 1: Change f(x) to y. Step 2: Flip x and y. So y will become x, and x will become y. Step 3: Make the function look like this--> y=mx+b

OpenStudy (anonymous):